# CLIPLoss and Norm-Based Data Selection Methods for Multimodal Contrastive Learning

### v1

@@ A.2 Generalization Guarantees
Theorem A.1 (Main).
Under the assumption of Lemma A.1,
@@
A.3 Detailed proofs
Lemma A.2.
@@ A.3 Detailed proofs
Note that $s_{ij}=({\bm{x}}_{j}^{l})^{\top}G_{l}G_{v}^{\top}{\bm{x}}_{i}^{v}=\operatorname{Tr}(G_{v}^{\top}{\bm{x}}_{i}^{v}({\bm{x}}_{j}^{l})^{\top}G_{l})$, like the proof of Corollary B.1. in [42], we have
@@
where $\bar{\bm{x}}^{vl}:=(\sum_{i\in S}{\bm{x}}_{i}^{vl})/|S|$. Then by the Eckart-Young-Mirsky Theorem (For example, Theorem 2.4.8 in Golub et al. [43]), we know that
@@
So $\frac{1}{|S|}\|\sum_{i}{\bm{z}}_{i}^{vl}(\xi_{i}^{vl})^{\top}\|_{*}\leq\sqrt{\frac{d\log(1/\delta)}{|S|}}$. Note that $\|\bar{\bm{x}}^{vl}\|\lesssim\sqrt{\frac{\log(|S|d)}{|S|}}$ (like Proposition 2.5 in Wainwright et al. [44]), it is easy to see that P3 ad P4 are the low order terms if $\delta\lesssim\frac{1}{|S|d}$. ∎
Lemma A.4 (Intuition behind VAS).
@@
Here the first equation comes from Theorem A.3 and the third equation comes from Lemma A.2.
@@
Theorem A.2 (Main).
Under the assumption of Lemma A.1, we have
@@
Based on Lemma A.1, we will focus on the error cause from selecting subset $S$, that is, $\operatorname{Tr}\Sigma_{\text{target}}\Sigma_{S}$. Since the exact $\Sigma_{\text{target}}$ is unknown, we assume the access to some proxy $\bar{\Sigma}_{\text{target}}$ instead.
@@
Theorem A.3 (A simplified version of test loss).
@@
Figure 4: Illustration of different directions for data selection methods for multimodal contrastive learning. Here we use four colors to denote the four main resources we can obtain: CLIP teacher model, downstream target data (which is much smaller than the external multimodal dataset or pretraining dataset), the external image-text dataset, and the external non-CLIP model. Direction 1 denotes the methods that only use the original OAI CLIP teacher model and the downstream target data. Direction 2 represents the methods that use external datasets to train a new CLIP teacher model for improving filtering, like DFN [2]. Direction 3 denotes the methods that use external non-CLIP model to select the data that may be heuristically helpful for downstream tasks, like image without too much text or be more special. In general, D1 method using only CLIP embedding, like negCLIPLoss, is orthogonal to D2. And both D1 and D2 can be combined with D3 to explore better filtering results. In the experiments part of the main paper (Sec. 4), we further show that our proposed D1 methods: NormSim and negCLIPLoss, can outperform all the D3 baselines except the best method “HYPE $\cup$ DFN”. And we can achieve the new state-of-the-art by combining our methods with that method.
@@
Appendix B Illstration of Different Directions for Data Selection for Multimodal Contrastive Learning
@@
We summarize our main idea of categorizing the current top data selection methods in Table 4.
Appendix C Details of Experiments
C.1 Computation Cost
@@
Table 4: Comparison of preprocessing time and external resources needed between our method and other D3 category methods. We skip DFN since it’s orthogonal to our negCLIPLoss method and we can directly improve it as mentioned in Table 2. Here since all the baselines below except MLM use a pretrained CLIP model, we only count the time that doesn’t contain that for inferring CLIP image/text embeddings (about 50 L40 hours for OAI CLIP-B/32), which is also adopted in DataComp benchmark [1]. The external dataset corresponds to the external multimodal dataset used for training or finetuning the external model. Notably, the preprocessing time for the following methods are all approximately linearly proportional to the amount of unfiltered pretrained dataset.
@@
D1 $\mathbb{D}^{2}$ Pruning [14] NA NA >70 L40 h 65 L40 h 29.5
D3 T-MARS [8] FAST [9] NA 950 L40 h 65 L40 h 34.1
D3 MLM [38] LLaVA-1.5 [39, 40] 50k 1120 A100 h 65 L40 h 34.5
D3 Devil [10] fasttext [11], BLIP-2 [12] NA 510 A100 h 65 L40 h 34.5
D3 HYPE [3] MERU [41] 27M > 120 L40 h 65 L40 h 31.9
@@
Our algorithm can significantly reduce the computational cost compared to many existing works as shown in Table 4. For example, when the CLIP embeddings are obtained (cost about 50 hours for CLIP-B/32), both T-MARS [8] and MLM [38] still require more than 900 hours data pre-processing time to extract the required information from 110M size dataset of DataComp-medium, while we only need about 5 hours. On the other hand, DFN, although has a similar forward speed (i.e. preprocessing time), requires retraining a new CLIP teacher model on the HQITP-350M, which is larger than DataComp-medium.
@@
  • For Devil, it needs to run the k-means clustering algorithm from the faiss library on the embedding space, which is estimated to cost 120 L40 hours on DataComp-medium. Using BLIP-2 [12] to scan the whole dataset will need about 470 A100 hours from the experimental details in [13]. From https://lambdalabs.com/gpu-benchmarks, we roughly assume that 120 L40 hours are at least comparable to 40 A100 hours for K-means clustering.
@@
C.2 Details of negCLIPLoss
In negCLIPLoss, we need to get the batch size $|B|$ and the value of the learnable temperature parameter $\tau$ at the final step of the teacher model pretraining stage. For OAI CLIP-L/14 and OAI CLIP-B/32, these values are $\tau=0.01$ and $|B|=32768$. We give the pseudocode of calculating negCLIPLoss in Algorithm 1, which is specially designed for pytorch to accelerate matrix calculation. It can be fully accelerated and the computation cost introduced by the normalization term is negligible compared with the training time or preprocessing time of other top baselines as detailed in Table C.1.
@@ C.2 Details of negCLIPLoss
C.4 More details in Baselines
@@
  • Text-based filtering. [1] proposes a text-based filtering that tries to select the data that contains caption overlapping with the class name from ImageNet-21K or ImageNet-1K.
  • Image-based filtering. [1] also proposes a heuristic way to sample the visual content overlaps with ImageNet-1K classes. They first apply filtering by language (only choose English caption by fasttext [11]) and caption length (over two words and 5 characters). Then they cluster the image embeddings from training data to 100K groups using Faiss [45], and keep the groups whose cluster center is the nearest neighbor to at least one image embedding of ImageNet-1K image.
  • $\mathbb{D}^{2}$ Pruning. [14] tries to represent the dataset as an undirected graph for coreset selection. They assign the difficulty for each example and use message passing to update the difficulty score incorporating the difficulty of its neighboring examples, and finally try to keep both diverse and difficult subsets. For our experiments, we adhere to the default hyperparameters of $\mathbb{D}^{2}$ on DataComp as specified in their official codebase.
  • T-MARS [8] uses a text detection model like FAST [9] to filter out the data that only contain the texts of caption in the image and don’t have other useful image features.
  • Devils [10] combines many ways for data filtering. At the very first it filter data based on heuristic rules like text length, frequency of texts, and image size, and it also use CLIPScore for cross-modality matchment. Then it adopts target distribution alignment methods similar to image-based filtering, but instead of using ImageNet-1k only, it uses 22 downstream tasks as the target set. Further, it adopts external models fasttext [11] to remove non-English captions and image-captioning model BLIP-2 [46] to select images with MNIST-style digits.
  • MLM [38] prompts GPT-4V to construct instruction data including the image-text data, and use it to fine-tune a smaller vision-language model like LLaVA-1.5 [39, 40] into a filtering network. Nevertheless, the number of parameters of LLaVA-1.5 is still much larger than CLIP, and thus LLaVA-1.5 has a much longer preprocessing time as mentioned in Table C.1.
C.5 Hyperparameters
The main hyper-parameters of our negCLIPLoss and NormSim are the target numbers for filtering, which is also the main concerns for all the top baselines like DFN, MLM, and T-MARS. However, from the current paper about the scaling law for data filtering [47], the downsampling size should depend on the computation budget. When you have more budget, you should sample more data for learning. And we can use their fitting formula to get some good downsampling ratios.
@@ C.5 Hyperparameters
Appendix D Additional Results
D.1 Stability Analysis of Batch Sampling Numbers in negCLIPLoss
@@
D.2 Universality of negCLIPLoss over Different Teacher Models
@@
OAI CLIP-B/32 Dataset IN-1k IN Dist. Shift VTAB Retrieval Avg.
Size (1 sub-task) (5) (11) (3) (38)
@@
D.3 $\text{NormSim}_{\infty}$ is Better than Nearest Neighbor Selection
@@
D.4 Vision-Only NormSim is Better than Using Both Vision and Language
@@
Appendix E Additional Visualization

### v2

@@ B.2 Generalization Guarantees
Theorem B.1 (Main).
Under the assumption of Lemma B.1,
@@
B.3 Detailed proofs
Lemma B.2.
@@ B.3 Detailed proofs
Note that $s_{ij}=({\bm{x}}_{j}^{l})^{\top}G_{l}G_{v}^{\top}{\bm{x}}_{i}^{v}=\operatorname{Tr}(G_{v}^{\top}{\bm{x}}_{i}^{v}({\bm{x}}_{j}^{l})^{\top}G_{l})$, like the proof of Corollary B.1. in (nakada2023understanding), we have
@@
where $\bar{\bm{x}}^{vl}:=(\sum_{i\in S}{\bm{x}}_{i}^{vl})/|S|$. Then by the Eckart-Young-Mirsky Theorem (For example, Theorem 2.4.8 in Golub et al. (golub2013matrix)), we know that
@@
So $\frac{1}{|S|}\|\sum_{i}{\bm{z}}_{i}^{vl}(\xi_{i}^{vl})^{\top}\|_{*}\leq\sqrt{\frac{d\log(1/\delta)}{|S|}}$. Note that $\|\bar{\bm{x}}^{vl}\|\lesssim\sqrt{\frac{\log(|S|d)}{|S|}}$ (like Proposition 2.5 in Wainwright et al. (wainwright2019high)), it is easy to see that P3 ad P4 are the low order terms if $\delta\lesssim\frac{1}{|S|d}$. ∎
Lemma B.4 (Intuition behind VAS).
@@
Here the first equation comes from Theorem B.3 and the third equation comes from Lemma B.2.
@@
Theorem B.2 (Main).
Under the assumption of Lemma B.1, we have
@@
Based on Lemma B.1, we will focus on the error cause from selecting subset $S$, that is, $\operatorname{Tr}\Sigma_{\text{target}}\Sigma_{S}$. Since the exact $\Sigma_{\text{target}}$ is unknown, we assume the access to some proxy $\bar{\Sigma}_{\text{target}}$ instead.
@@
Theorem B.3 (A simplified version of test loss).
@@
Figure 4: Illustration of different directions for data selection methods for multimodal contrastive learning. Here we use four colors to denote the four main resources we can obtain: CLIP teacher model, downstream target data (which is much smaller than the external multimodal dataset or pretraining dataset), the external image-text dataset, and the external non-CLIP model. Direction 1 denotes the methods that only use the original OAI CLIP teacher model and the downstream target data. Direction 2 represents the methods that use external datasets to train a new CLIP teacher model for improving filtering, like DFN (fang2023data). Direction 3 denotes the methods that use external non-CLIP model to select the data that may be heuristically helpful for downstream tasks, like image without too much text or be more special. In general, D1 method using only CLIP embedding, like negCLIPLoss, is orthogonal to D2. And both D1 and D2 can be combined with D3 to explore better filtering results. In the experiments part of the main paper (Sec. 4), we further show that our proposed D1 methods: NormSim and negCLIPLoss, can outperform all the D3 baselines except the best method “HYPE $\cup$ DFN”. And we can achieve the new state-of-the-art by combining our methods with that method.
@@
Appendix C Illstration of Different Directions for Data Selection for Multimodal Contrastive Learning
@@
We summarize our main idea of categorizing the current top data selection methods in Table 4. In our experiment parts, we consider the baselines below:
D1: OAI CLIP embedding only.
The learner can only access the pretraining dataset (like DataComp-medium), the original OAI CLIP teacher model that is used to extract embeddings, and some target data of the downstream tasks which is much smaller than the pretraining dataset (like ImageNet-1k). In this category, we don’t use any existing external non-CLIP models or any newly trained CLIP model based on external multi-modal dataset. In detail, This category includes (1) CLIPScore (hessel2021clipscore), which only uses CLIPScore for filtering as we mentioned before. (2) Image-based filtering (gadre2023datacomp), which uses ImageNet-1K training data as the downstream target data for data filtering. It applies k-means clustering to the image embeddings of training data and selects clusters closest to the ImageNet-1K embeddings. Gadre et al. (gadre2023datacomp) also try to combine image-based filtering and CLIPScore together. (3) $\mathbb{D}^{2}$ Pruning (maharana2023d2), which represents the dataset as an undirected graph and selects the data by combining difficulty and diversity. They use the CLIP score to initialize their graph.
D2, D3: Accessible external model and multi-modal data.
All the current top baselines enable the learner to utilize external resources, either to train a better CLIP teacher model or to help filtering using existing models’ properies. In detail, (1) DFN (fang2023data) trains another CLIP data filtering network via external high-quality data. Their currently public model (DFN-P) is trained on CC12M (changpinyo2021conceptual12) + CC3M (sharma2018conceptual3) + SS15M (nguyen2022quality), while the best DFN is trained on nonpublic HQITP-350M (fang2023data), which is even larger than DataComp-medium. (2) HYPE (kim2024hype) leverages hyperbolic embeddings (different from CLIP embedding) and the concept of entailment cones to filter out samples with meaningless or underspecified semantics, enhancing the specificity of each sample. (3) HYPE $\cup$ DFN proposed by (kim2024hype) samples subset separately for each method and then merge them. This is the state-of-the-art method on the DataComp benchmark for medium size. (4) Other methods including T-MARS (maini2023t), Devils (yu2023devil), MLM (wang2024finetuned), which leverage external models such as text detection model FAST (chen2021fast), vision-language model (VLM) BLIP-2 (li2023blip) and VLM LLaVA-1.5 (liu2023improved; chiang2023vicuna) to heuristically select the more helpful data, like non-text-only pictures or images with MNIST-style digits. See details in Appendix D.5.
Appendix D Details of Experiments
D.1 Details of Setup
Evaluation.
We measured the model performance on 38 downstream datasets including image classification and retrieval tasks followed by DataComp. The image classification tasks contain ImageNet-1k (deng2009imagenet), ImageNet distribution shifts (wang2019learning; pmlr-v97-recht19a; hendrycks2021natural; hendrycks2021many), 11 datasets from the Visual Task Adaptation Benchmark (VTAB) (zhai2019large) and 3 datasets from WILDS (koh2021wilds; sagawa2021extending). Retrieval datasets contain Flickr30k (young-etal-2014-image), MSCOCO (chen2015microsoft) and WinoGAViL (bitton2022winogavil).
Teacher model architecture.
Our experiments utilize two architectures for OpenAI’s CLIP teacher models: ViT-L/14 and ViT-B/32. Additionally, we use the public version of DFN (DFN-P) proposed by Fang et al. (fang2023data) as a teacher model, and its architecture is also ViT-B/32.
D.2 Computation Cost
@@
Table 4: Comparison of preprocessing time and external resources needed between our method and other D3 category methods. We skip DFN since it’s orthogonal to our negCLIPLoss method and we can directly improve it as mentioned in Table 2. Here since all the baselines below except MLM use a pretrained CLIP model, we only count the time that doesn’t contain that for inferring CLIP image/text embeddings (about 50 L40 hours for OAI CLIP-B/32), which is also adopted in DataComp benchmark (gadre2023datacomp). The external dataset corresponds to the external multimodal dataset used for training or finetuning the external model. Notably, the preprocessing time for the following methods are all approximately linearly proportional to the amount of unfiltered pretrained dataset.
@@
D1 $\mathbb{D}^{2}$ Pruning NA NA >70 L40 h 65 L40 h 29.5
D3 T-MARS FAST NA 950 L40 h 65 L40 h 34.1
D3 MLM LLaVA-1.5 50k 1120 A100 h 65 L40 h 34.5
D3 Devil fasttext, BLIP-2 NA 510 A100 h 65 L40 h 34.5
D3 HYPE MERU 27M > 120 L40 h 65 L40 h 31.9
@@
Our algorithm can significantly reduce the computational cost compared to many existing works as shown in Table 4. For example, when the CLIP embeddings are obtained (cost about 50 hours for CLIP-B/32), both T-MARS (maini2023t) and MLM (wang2024finetuned) still require more than 900 hours data pre-processing time to extract the required information from 110M size dataset of DataComp-medium, while we only need about 5 hours. On the other hand, DFN, although has a similar forward speed (i.e. preprocessing time), requires retraining a new CLIP teacher model on the HQITP-350M, which is larger than DataComp-medium.
@@
  • For Devil, it needs to run the k-means clustering algorithm from the faiss library on the embedding space, which is estimated to cost 120 L40 hours on DataComp-medium. Using BLIP-2 (li2023blip) to scan the whole dataset will need about 470 A100 hours from the experimental details in (nguyen2023improving). From https://lambdalabs.com/gpu-benchmarks, we roughly assume that 120 L40 hours are at least comparable to 40 A100 hours for K-means clustering.
@@
D.3 Details of negCLIPLoss
In negCLIPLoss, we need to get the batch size $|B|$ and the value of the learnable temperature parameter $\tau$ at the final step of the teacher model pretraining stage. For OAI CLIP-L/14 and OAI CLIP-B/32, these values are $\tau=0.01$ and $|B|=32768$. We give the pseudocode of calculating negCLIPLoss in Algorithm 1, which is specially designed for pytorch to accelerate matrix calculation. It can be fully accelerated and the computation cost introduced by the normalization term is negligible compared with the training time or preprocessing time of other top baselines as detailed in Table D.2.
@@ D.3 Details of negCLIPLoss
D.5 More details in Baselines
@@
  • Text-based filtering. (gadre2023datacomp) proposes a text-based filtering that tries to select the data that contains caption overlapping with the class name from ImageNet-21K or ImageNet-1K.
  • Image-based filtering. (gadre2023datacomp) also proposes a heuristic way to sample the visual content overlaps with ImageNet-1K classes. They first apply filtering by language (only choose English caption by fasttext (joulin2016bag)) and caption length (over two words and 5 characters). Then they cluster the image embeddings from training data to 100K groups using Faiss (johnson2019billion), and keep the groups whose cluster center is the nearest neighbor to at least one image embedding of ImageNet-1K image.
  • $\mathbb{D}^{2}$ Pruning. (maharana2023d2) tries to represent the dataset as an undirected graph for coreset selection. They assign the difficulty for each example and use message passing to update the difficulty score incorporating the difficulty of its neighboring examples, and finally try to keep both diverse and difficult subsets. For our experiments, we adhere to the default hyperparameters of $\mathbb{D}^{2}$ on DataComp as specified in their official codebase.
  • T-MARS (maini2023t) uses a text detection model like FAST (chen2021fast) to filter out the data that only contain the texts of caption in the image and don’t have other useful image features.
  • Devils (yu2023devil) combines many ways for data filtering. At the very first it filter data based on heuristic rules like text length, frequency of texts, and image size, and it also use CLIPScore for cross-modality matchment. Then it adopts target distribution alignment methods similar to image-based filtering, but instead of using ImageNet-1k only, it uses 22 downstream tasks as the target set. Further, it adopts external models fasttext (joulin2016bag) to remove non-English captions and image-captioning model BLIP-2 (nguyen2024improving) to select images with MNIST-style digits.
  • MLM (wang2024finetuned) prompts GPT-4V to construct instruction data including the image-text data, and use it to fine-tune a smaller vision-language model like LLaVA-1.5 (liu2023improved; chiang2023vicuna) into a filtering network. Nevertheless, the number of parameters of LLaVA-1.5 is still much larger than CLIP, and thus LLaVA-1.5 has a much longer preprocessing time as mentioned in Table D.2.
D.6 Hyperparameters
The main hyper-parameters of our negCLIPLoss and NormSim are the target numbers for filtering, which is also the main concerns for all the top baselines like DFN, MLM, and T-MARS. However, from the current paper about the scaling law for data filtering (goyal2024scaling), the downsampling size should depend on the computation budget. When you have more budget, you should sample more data for learning. And we can use their fitting formula to get some good downsampling ratios.
@@ D.6 Hyperparameters
Appendix E Additional Results
E.1 Stability Analysis of Batch Sampling Numbers in negCLIPLoss
@@
E.2 Complete Comparison Results for D1 categories
[…]
Filtering Strategy Dataset IN-1k IN Shift VTAB Retr. Avg.
Size (1 task) (5) (11) (3) (38)
No filtering (gadre2023datacomp) 110M 17.3 15.0 25.2 21.3 25.6
CLIPScore (20%) (hessel2021clipscore) 22M 25.4 22.7 31.8 22.0 31.0
CLIPScore (30%) (hessel2021clipscore) 33M 26.4 23.6 32.6 24.5 32.2
Image-based (gadre2023datacomp) 24M 25.5 21.9 30.4 24.6 29.9
CLIPScore (30%) $\cap$ Image-based 11M 27.4 23.9 31.9 21.4 30.8
$\mathbb{D}^{2}$ Pruning (maharana2023d2) 22M 23.2 20.4 31.4 18.7 29.5
negCLIPLoss (20%) 22M 27.4 23.8 33.7 23.7 32.5
negCLIPLoss (30%) 33M 27.9 24.6 33.2 25.1 32.9
[…]
E.3 Universality of negCLIPLoss over Different Teacher Models
@@
OAI CLIP-B/32 Dataset IN-1k IN Shift VTAB Retr. Avg.
Size (1) (5) (11) (3) (38)
@@
E.4 $\text{NormSim}_{\infty}$ is Better than Nearest Neighbor Selection
@@
E.5 Vision-Only NormSim is Better than Using Both Vision and Language
@@
Appendix F Additional Visualization